OpenStudy (anonymous):

In the values interval [0, 2p] the function f (x) = sin(x) has more periods by the function g (x) = sin(3x). True or False?

OpenStudy (anonymous):

The period of \(\sin bx\) is \(\dfrac{2\pi}{b}\). Is the period of \(f(x)\) greater than the period of \(g(x)\) ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

@SithsAndGiggles because you subtract with 1 (x) instead of 3 to the second one (3x).. So the period of f(x) is greater than the period of g(x) right?

OpenStudy (anonymous):

Well, the period of \(f\) is \(2\pi\), whereas the period of \(g\) is \(\dfrac{2\pi}{3}\). I'm not sure about your reasoning, but it's clear that you know \(f\)'s period is greater.